OpenStudy (anonymous):

Can someone help me with exponents? (8th grade level)

OpenStudy (anonymous):

What help do you need?

OpenStudy (anonymous):

@mrthompson44

OpenStudy (anonymous):

@mrthompson440

OpenStudy (anonymous):

@mthompson440

OpenStudy (anonymous):

To multiply the exponents you add them. To divide them you subtract. You first need to add the exponents in the numerator together, then you will subtract the exponent in the denominator from the numerator.

OpenStudy (anonymous):

For example: A^4 * A^3= A^7 A^7/A^2= A^5